OpenStudy (anonymous):

A regular heptagonal prism has a volume of 52m^3. If each one of its edges is doubled in size, what would be its new volume? I know the answer its 416m^3 but I don't know how to solve it. Thanks!! :)

OpenStudy (jack1):

formula for volume of a prism = area of base x ht so what is happening to the base area if you double one of the sides?

OpenStudy (anonymous):

it increases

OpenStudy (jack1):

yeah... but by a factor of...? ie whats the formula for area of a heptagon?

OpenStudy (anonymous):

its 5 x apothem x perimeter?

OpenStudy (jack1):

yeah, that's one formula but in this case i'd go with the regular polygon formula Area = (1/4 x s^2 x N) / (tan [180/N])

OpenStudy (anonymous):

but I don't get how do you know by what factor does it increases is it 2 because it says that it double the sides or do you count the sides? How do you know?

OpenStudy (jack1):

now it looks complex, but it's really simple and it works for all regular polygons the key is s = side length; and N = number of sides so for a heptagon: 7 sides Area = (1/4 x s^2 x 7) / (tan [180/7]) Area = (7/4 x s^2 ) / (tan [180/25.71]) Area = (1.75 x s^2) / (0.657) therefore Area of heptagon = 2.66 * s^2

OpenStudy (jack1):

so assume a side length of 10 and ht of 10 for a sec: Area of base = 2.66 x 10^2 = 2.66 x 100 ht remains the same, so volume = 2660 now if you double side length: new side length of 2x10 so new Area of base = 2.66 x (2x10)^2 = 2.66 x 20^2 = 2.66 x 400 = 1064 and ht of 10 remains the same, so volume = 10640 as u can see, increasing side length by 2times increased volume by 4times (2^2) so volume varies with the square of side length

OpenStudy (jack1):

ie if you increased side length to 4 times the original size, the change in volume would be 16 times greater (4x4 = 16)

OpenStudy (jack1):

from your question it states "if each one of the edges doubles in size", therefore side length doubles

OpenStudy (anonymous):

oh I get it so if it 3 times the original size the change in volume would be 9 (3X3=9)

OpenStudy (jack1):

yep, u got it so volume varies with (side length) sqrd

OpenStudy (jack1):

Mathematically from Q: so original volume of 52: (from q) assume height of 4 original side length would be Area x ht = volume 2.66 x s^2 x 4 = 52 2.66 x s^2 = 13 s^2 = 4.8872180451 s = 2.2107 now if you double the side length of the heptagon: s = 2.2107 x 2 = 4.4114 Volume = area of base x ht = 2.66 x s^2 x 4 = 2.66 x (4.4114)^2 x 4 = 207.01 ish which is roughly 208 which is 4 times the volume of the original

OpenStudy (anonymous):

ok

OpenStudy (jack1):

aw damm, ëach one of its edges is doubled in size so the height is also doubled so : = 2.66 x s^2 x 8 = 2.66 x (4.4114)^2 x 8 = 414.11 ish which is roughly 416 (as i've been rounding) which is 8 times the volume of the original

OpenStudy (jack1):

so summation: double height = 2x volume double side length of the base = 4x volume do both = 4x2x volume = 8x volume

OpenStudy (anonymous):

oh ok it make sense now! Thank you so much now I think I get it, thank you I really didn't know how to solve it

OpenStudy (jack1):

all good, sorry it took so long dude

OpenStudy (anonymous):

It's ok I didn't mind waiting but thank you for helping me :)

OpenStudy (jack1):

nah any time dude, good luck and slaters ;D

OpenStudy (anonymous):

thanks! later :D

Directrix (directrix):

Theorem: If two solids are similar, the cube of the scale factor of the two solids is equal to the ratio of the volumes. (1/2)^3 = 52/x where x is the new volume and 1/2 is the scale factor. 1/8 = 52/x x = 8 * 52 = ?

OpenStudy (anonymous):

thanks Directrix :)

Directrix (directrix):

You are welcome.